OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (campbell_st):

just swap x and y, then make y the subject of the equation so swapping you have \[x = y^2 - 16\] now make y the subject

OpenStudy (jdoe0001):

\(\bf f(x)= x^2 - 16 \implies \color{red}{y} = \color{blue}{x}^2 - 16\\ f^{-1}(x)\implies \color{blue}{x} = \color{red}{y}^2 - 16\)